i need help to solve this by the substitution method please, can you please show me steps to how you solve this, thank you



2x-y=5 9x-4y=23

Answer by ankor@dixie-net.com(22740)

2x - y = 5
9x- 4y = 23
:
Use the 1st equation for substitution
2x - y = 5
Subtract 2x from both sides
-y = -2x + 5
y has to be positive, multiply both sides by -1
y = 2x - 5
:
In the 2nd equation,
9x - 4y = 23
substitute (2x-5) for y
9x - 4(2x-5) = 23
9x - 8x + 20 = 23
subtract 20 from both sides
x = 23 - 20
x = 3
:
We know y = 2x -5
substitute x with 3
y = 2(3) - 5
y = 6 - 5
y = 1
:
Our solution x=3, y=1 should be checked in both equations
2x - y = 5
replace x and y
2(3) - 1 = 5
and
9x- 4y = 23
9(3) - 4(1) = 23
:
HOw about this, enough steps for you to understand what happened here? C